# Continuous-Time Quantum Markov Chains And Discretizations Of $p$-Adic Schrödinger Equations: Comparisons And Simulations

### v1

@@ 11. Appendix D: $p$-adic Schrödinger equations on the unit ball and quantum Markov chains
(11.1) $\left\{\begin{array}[c]{l}i\frac{\partial}{\partial t}\Psi\left(x,t\right)=-J\left(\left|x\right|_{p}\right)\ast\Psi\left(x,t\right)+\Psi\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t\geq 0\\ \\ \Psi\left(x,0\right)=\psi_{0}\left(x\right)\in\mathcal{D}(\mathbb{Z}_{p}).\end{array}\right$.
@@
(11.2) $\Psi\left(x,t\right)=\mathcal{F}_{\xi\rightarrow x}^{-1}(e^{-i\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t})\ast\psi_{0}\left(x\right)\text{ in }\mathcal{D}^{\prime}(\mathbb{Q}_{p})\text{,}$
@@
$e^{-i\left(1-\widehat{J}\left(\left|\xi+\xi_{0}\right|_{p}\right)\right)t}=e^{-i\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t}$
for any $\xi_{0}\in\mathbb{Z}_{p}$, $\xi\in\mathbb{Q}_{p}$, $\mathcal{F}_{\xi\rightarrow x}^{-1}(e^{-i\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t})$ is a distribution with support in $\mathbb{Z}_{p}$, see, e.g., [19, Section VII.3], [35, Theorem 4.9.3], [32, Theorem 1.161]. The explicit calculation of $\mathcal{F}_{\xi\rightarrow x}^{-1}(e^{-i\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t})$ can be carried out as in the case of the heat kernel, and one obtains that
$\mathcal{F}_{\xi\rightarrow x}^{-1}(e^{-i\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t}))=Z_{0}(x,it)$
@@
$=\Omega\left(\left|x\right|_{p}\right){\displaystyle\int\limits_{\mathbb{Z}_{p}}}J\left(\left|x-y\right|_{p}\right)\Omega\left(\left|y\right|_{p}\right)dy-\Omega\left(\left|x\right|_{p}\right)=\Omega\left(\left|x\right|_{p}\right){\displaystyle\int\limits_{\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy-\Omega\left(\left|x\right|_{p}\right)$
$=\Omega\left(\left|x\right|_{p}\right){\displaystyle\int\limits_{\mathbb{Z}_{p}}}J\left(\left|y\right|_{p}\right)dy-\Omega\left(\left|x\right|_{p}\right)=\Omega\left(\left|x\right|_{p}\right)\left(\left\|J\right\|_{L^{1}\left(\mathbb{Z}_{p}\right)}-1\right).$
@@
(11.4) $\Psi(x,t)=A_{0}\Omega\left(\left|x\right|_{p}\right)+{\displaystyle\sum\limits_{k\in\left\{1,\ldots,p-1\right\}}}\text{ \ }{\displaystyle\sum\limits_{\begin{subarray}{c}b\in\mathbb{Q}_{p}/\mathbb{Z}_{p}\\ bp^{-r}\in\mathbb{Z}_{p}\end{subarray}}}\text{ \ }{\displaystyle\sum\limits_{r\leq 0}}\text{ }A_{{}_{rbk}}e^{-i\left(J\left(p^{1-r}\right)-1\right)t}\Psi_{rbk}\left(x\right){,}$
@@ 11.1. CTQMCs
We use all the notation introduced in Subsection $3.$ Set $e_{v}=p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)$ for $v\in G_{l}$. Then $\left\{p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\right\}_{v\in G_{l}}$ is an orthonormal basis of the finite-dimensional Hilbert space $\mathcal{H}_{G_{l}}=\mathcal{D}_{l}(\mathbb{Z}_{p})$ Furthermore,
$e^{it\boldsymbol{J}}p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)=Z_{0}(x,it)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\in\mathcal{D}_{l}(\mathbb{Z}_{p}),$
cf. Lemma 9.5-(iii). This fact implies that $e^{it\boldsymbol{J}}\mathcal{D}_{l}(\mathbb{Z}_{p})\subset\mathcal{D}_{l}(\mathbb{Z}_{p})$. Now, the announced result follows from Theorem 3.1 . ∎
@@
$\alpha_{r,v}\left(t\right):=\left\langle p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right),Z_{0}(x,it)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\right\rangle,$
@@
(11.5) $\alpha_{r,v}\left(t\right)=p^{-l}\left\{e^{it}{\displaystyle\int\limits_{\left|\xi\right|_{p}\leq p^{l}}}\chi_{p}\left(\left(r-v\right)\xi\right)e^{-it\widehat{J}(\left|\xi\right|_{p})}d\xi\right\},$
@@
$Z_{0}(x,it)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\in L^{2}\left(\mathbb{Z}_{p}\right)\subset L^{2}\left(\mathbb{Q}_{p}\right).$
Using that the Fourier transform preserves the inner product in $L^{2}\left(\mathbb{Q}_{p}\right)$,  we have
$\left\langle p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right),Z_{0}(x,it)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\right\rangle$
$=p^{-l}\left\langle\Omega\left(p^{-l}\left|\xi\right|_{p}\right)\chi_{p}\left(r\xi\right),e^{it\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\Omega\left(p^{-l}\left|\xi\right|_{p}\right)\chi_{p}\left(v\xi\right)\right\rangle$
$=p^{-l}{\displaystyle\int\limits_{\mathbb{Q}_{p}}}\Omega\left(p^{-l}\left|\xi\right|_{p}\right)\chi_{p}\left(r\xi\right)e^{-it\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\Omega\left(p^{-l}\left|\xi\right|_{p}\right)\chi_{p}\left(-v\xi\right)d\xi$
$=p^{-l}{\displaystyle\int\limits_{\left|\xi\right|_{p}\leq p^{l}}}\chi_{p}\left(\left(r-v\right)\xi\right)e^{-it\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}d\xi.$
@@
$\left[\pi_{r,v}\left(t\right)\right]_{r,v\in G_{l}}=\left[p^{-2l}\left|\text{ }{\displaystyle\int\limits_{\left|\xi\right|_{p}\leq p^{l}}}\chi_{p}\left(\left(r-v\right)\xi\right)e^{-it\widehat{J}(\left|\xi\right|_{p})}d\xi\right|^{2}\right]_{r,v\in G_{l}}$
@@ 12.2. Discretizations of differential equations
(12.6) $\left\{\begin{array}[c]{l}\left[\Psi_{I}^{\left(l\right)}\left(t\right)\right]\in\mathbb{C}^{p^{l}}\text{, }t\geq 0;\text{ }\Psi_{I}^{\left(l\right)}\left(t\right)\in C^{1}(\mathbb{R}_{+})\text{, }I\in G_{l}\\ \\ i\frac{\partial}{\partial t}\left[\Psi_{I}^{\left(l\right)}\left(t\right)\right]=-\boldsymbol{J}^{\left(l\right)}\left[\Psi_{I}^{\left(l\right)}\left(t\right)\right]\text{, }I\in G_{l},t\geq 0\\ \\ \left[\Psi_{I}^{\left(l\right)}\left(0\right)\right]=\left[\Psi_{I}^{\left(\text{init}\right)}\right]\in\mathbb{C}^{p^{l}},\end{array}\right$.
@@ 12.3. Quantum networks
(12.7) $\pi_{I,J}\left(t\right)=\left|\left\langle e_{I}\right|e^{it\boldsymbol{J}^{\left(l\right)}}\left|e_{J}\right\rangle_{l}\right|^{2}\text{, for }J,I\in G_{l}.$
@@
${\displaystyle\sum\limits_{I\in G_{l}}}\pi_{I,J}\left(t\right)={\displaystyle\sum\limits_{I\in G_{l}}}\left|\left\langle e_{I}\right|e^{it\boldsymbol{J}^{\left(l\right)}}\left|e_{J}\right\rangle\right|^{2}=1,$
@@ 12.4. Additional remarks
$\boldsymbol{P}_{l}e^{it\boldsymbol{J}}:\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)\rightarrow\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right),$
@@
$e^{it\boldsymbol{J}^{\left(l\right)}}:\mathcal{H}_{l}\rightarrow\mathcal{H}_{l}.$
@@
$\boldsymbol{P}_{l}e^{it\boldsymbol{J}}=e^{it\boldsymbol{P}_{l}\boldsymbol{J}}=e^{it\boldsymbol{J}^{\left(l\right)}}.$
@@
$e^{it\boldsymbol{P}_{l}\boldsymbol{J}}={\displaystyle\sum\limits_{k=0}^{\infty}}\frac{\left(it\right)^{k}}{k!}\left(\boldsymbol{P}_{l}\boldsymbol{J}\right)^{k}={\displaystyle\sum\limits_{k=0}^{\infty}}\frac{\left(it\right)^{k}}{k!}\boldsymbol{P}_{l}\boldsymbol{J}^{k}=\boldsymbol{P}_{l}{\displaystyle\sum\limits_{k=0}^{\infty}}\frac{\left(it\right)^{k}}{k!}\boldsymbol{J}^{k}=\boldsymbol{P}_{l}e^{it\boldsymbol{J}}.$
@@
$\boldsymbol{P}_{l}e^{t\boldsymbol{J}}=e^{t\boldsymbol{P}_{l}\boldsymbol{J}}=e^{t\boldsymbol{J}^{\left(l\right)}}:\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)\rightarrow\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right).$
@@
$p_{r,v}(t)$ $=p^{l}{\displaystyle\int\limits_{v+p^{l}\mathbb{Z}_{p}}}Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)dx=p^{l}{\displaystyle\int\limits_{v+p^{l}\mathbb{Z}_{p}}}e^{t\boldsymbol{J}^{\left(l\right)}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)dx$
$={\displaystyle\int\limits_{\mathbb{Z}_{p}}}p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)e^{t\boldsymbol{J}^{\left(l\right)}}p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)dx,$

### v2

@@ 11. Appendix D: $p$-adic Schrödinger equations on the unit ball and quantum Markov chains
(11.1) $\left\{\begin{array}[c]{l}\mathrm{i}\frac{\partial}{\partial t}\Psi\left(x,t\right)=-J\left(\left|x\right|_{p}\right)\ast\Psi\left(x,t\right)+\Psi\left(x,t\right)\text{, }x\in\mathbb{Z}_{p},t\geq 0\\ \\ \Psi\left(x,0\right)=\psi_{0}\left(x\right)\in\mathcal{D}(\mathbb{Z}_{p}).\end{array}\right$.
@@
(11.2) $\Psi\left(x,t\right)=\mathcal{F}_{\xi\rightarrow x}^{-1}(\mathrm{e}^{-\mathrm{i}\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t})\ast\psi_{0}\left(x\right)\text{ in }\mathcal{D}^{\prime}(\mathbb{Q}_{p})\text{,}$
@@
$\mathrm{e}^{-i\left(1-\widehat{J}\left(\left|\xi+\xi_{0}\right|_{p}\right)\right)t}=\mathrm{e}^{-i\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t}$
for any $\xi_{0}\in\mathbb{Z}_{p}$, $\xi\in\mathbb{Q}_{p}$, $\mathcal{F}_{\xi\rightarrow x}^{-1}(\mathrm{e}^{-\mathrm{i}\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t})$ is a distribution with support in $\mathbb{Z}_{p}$, see, e.g., [19, Section VII.3], [35, Theorem 4.9.3], [32, Theorem 1.161]. The explicit calculation of $\mathcal{F}_{\xi\rightarrow x}^{-1}(\mathrm{e}^{-\mathrm{i}\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t})$ can be carried out as in the case of the heat kernel, and one obtains that
$\mathcal{F}_{\xi\rightarrow x}^{-1}(\mathrm{e}^{-\mathrm{i}\left(1-\widehat{J}\left(\left|\xi\right|_{p}\right)\right)t}))=Z_{0}(x,\mathrm{i}t)$
@@
$=\Omega\left(\left|x\right|_{p}\right)\int\limits_{\mathbb{Z}_{p}}J\left(\left|x-y\right|_{p}\right)\Omega\left(\left|y\right|_{p}\right)dy-\Omega\left(\left|x\right|_{p}\right)$
$=\Omega\left(\left|x\right|_{p}\right)\int\limits_{\mathbb{Z}_{p}}J\left(\left|y\right|_{p}\right)dy-\Omega\left(\left|x\right|_{p}\right)$
$=\Omega\left(\left|x\right|_{p}\right)\int\limits_{\mathbb{Z}_{p}}J\left(\left|y\right|_{p}\right)dy-\Omega\left(\left|x\right|_{p}\right)=\Omega\left(\left|x\right|_{p}\right)\left(\left\|J\right\|_{L^{1}\left(\mathbb{Z}_{p}\right)}-1\right).$
@@
(11.4) $\Psi(x,t)=A_{0}\Omega\left(\left|x\right|_{p}\right)+{\displaystyle\sum\limits_{k\in\left\{1,\ldots,p-1\right\}}}\text{ \ }{\displaystyle\sum\limits_{\begin{subarray}{c}b\in\mathbb{Q}_{p}/\mathbb{Z}_{p}\\ bp^{-r}\in\mathbb{Z}_{p}\end{subarray}}}\text{ \ }{\displaystyle\sum\limits_{r\leq 0}}\text{ }A_{{}_{rbk}}\mathrm{e}^{-\mathrm{i}\left(J\left(p^{1-r}\right)-1\right)t}\Psi_{rbk}\left(x\right){,}$
@@ 11.1. CTQMCs
$\pi_{r,v}\left(t\right)=\left|\left\langle p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right),Z_{0}(x,\mathrm{i}t)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\right\rangle\right|^{2},$
@@
We use all the notation introduced in Subsection $3.$ Set
$e_{v}=p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right),$
for $v\in G_{l}$. Then $\left\{p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\right\}_{v\in G_{l}}$ is an orthonormal basis of the finite-dimensional Hilbert space $\mathcal{H}_{G_{l}}=\mathcal{D}_{l}(\mathbb{Z}_{p})$ Furthermore,
$\mathrm{e}^{\mathrm{i}t\boldsymbol{J}}p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)=Z_{0}(x,\mathrm{i}t)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\in\mathcal{D}_{l}(\mathbb{Z}_{p}),$
cf. Lemma 9.5-(iii). This fact implies that $\mathrm{e}^{\mathrm{i}t\boldsymbol{J}}\mathcal{D}_{l}(\mathbb{Z}_{p})\subset\mathcal{D}_{l}(\mathbb{Z}_{p})$. Now, the announced result follows from Theorem 3.1. ∎
@@
$\alpha_{r,v}\left(t\right):=\left\langle p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right),Z_{0}(x,\mathrm{i}t)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\right\rangle,$
@@
(11.5) $\alpha_{r,v}\left(t\right)=p^{-l}\left\{\mathrm{e}^{\mathrm{i}t}{\displaystyle\int\limits_{\left|\xi\right|_{p}\leq p^{l}}}\chi_{p}\left(\left(r-v\right)\xi\right)e^{-it\widehat{J}(\left|\xi\right|_{p})}d\xi\right\},$
@@
$Z_{0}(x,\mathrm{i}t)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\in L^{2}\left(\mathbb{Z}_{p}\right)\subset L^{2}\left(\mathbb{Q}_{p}\right).$
Using the fact that the Fourier transform preserves the inner product in $L^{2}\left(\mathbb{Q}_{p}\right)$, we have
$\left\langle p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right),Z_{0}(x,\mathrm{i}t)\ast p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\right\rangle$
$=p^{-l}\left\langle\Omega\left(p^{-l}\left|\xi\right|_{p}\right)\chi_{p}\left(r\xi\right),\mathrm{e}^{\mathrm{i}t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\Omega\left(p^{-l}\left|\xi\right|_{p}\right)\chi_{p}\left(v\xi\right)\right\rangle$
$=p^{-l}{\displaystyle\int\limits_{\mathbb{Q}_{p}}}\Omega\left(p^{-l}\left|\xi\right|_{p}\right)\chi_{p}\left(r\xi\right)\mathrm{e}^{-\mathrm{i}t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}\Omega\left(p^{-l}\left|\xi\right|_{p}\right)\chi_{p}\left(-v\xi\right)d\xi$
$=p^{-l}{\displaystyle\int\limits_{\left|\xi\right|_{p}\leq p^{l}}}\chi_{p}\left(\left(r-v\right)\xi\right)\mathrm{e}^{-\mathrm{i}t\left(\widehat{J}(\left|\xi\right|_{p})-1\right)}d\xi.$
@@
$\left[\pi_{r,v}\left(t\right)\right]_{r,v\in G_{l}}=\left[p^{-2l}\left|\text{ }{\displaystyle\int\limits_{\left|\xi\right|_{p}\leq p^{l}}}\chi_{p}\left(\left(r-v\right)\xi\right)\mathrm{e}^{-\mathrm{i}t\widehat{J}(\left|\xi\right|_{p})}d\xi\right|^{2}\right]_{r,v\in G_{l}}$
@@ 12.2. Discretizations of differential equations
(12.6) $\left\{\begin{array}[c]{l}\left[\Psi_{I}^{\left(l\right)}\left(t\right)\right]\in\mathbb{C}^{p^{l}}\text{, }t\geq 0;\text{ }\Psi_{I}^{\left(l\right)}\left(t\right)\in C^{1}(\mathbb{R}_{+})\text{, }I\in G_{l}\\ \\ \mathrm{i}\frac{\partial}{\partial t}\left[\Psi_{I}^{\left(l\right)}\left(t\right)\right]=-\boldsymbol{J}^{\left(l\right)}\left[\Psi_{I}^{\left(l\right)}\left(t\right)\right]\text{, }I\in G_{l},t\geq 0\\ \\ \left[\Psi_{I}^{\left(l\right)}\left(0\right)\right]=\left[\Psi_{I}^{\left(\text{init}\right)}\right]\in\mathbb{C}^{p^{l}},\end{array}\right$.
@@ 12.3. Quantum networks
(12.7) $\pi_{I,J}\left(t\right)=\left|\left\langle e_{I}\right|\mathrm{e}^{\mathrm{i}t\boldsymbol{J}^{\left(l\right)}}\left|e_{J}\right\rangle_{l}\right|^{2}\text{, for }J,I\in G_{l}.$
@@
${\displaystyle\sum\limits_{I\in G_{l}}}\pi_{I,J}\left(t\right)={\displaystyle\sum\limits_{I\in G_{l}}}\left|\left\langle e_{I}\right|\mathrm{e}^{\mathrm{i}t\boldsymbol{J}^{\left(l\right)}}\left|e_{J}\right\rangle\right|^{2}=1,$
@@ 12.4. Additional remarks
$\boldsymbol{P}_{l}\mathrm{e}^{\mathrm{i}t\boldsymbol{J}}:\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)\rightarrow\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right),$
@@
$\mathrm{e}^{\mathrm{i}t\boldsymbol{J}^{\left(l\right)}}:\mathcal{H}_{l}\rightarrow\mathcal{H}_{l}.$
@@
$\boldsymbol{P}_{l}\mathrm{e}^{\mathrm{i}t\boldsymbol{J}}=\mathrm{e}^{\mathrm{i}t\boldsymbol{P}_{l}\boldsymbol{J}}=\mathrm{e}^{\mathrm{i}t\boldsymbol{J}^{\left(l\right)}}.$
@@
$\mathrm{e}^{\mathrm{i}t\boldsymbol{P}_{l}\boldsymbol{J}}={\displaystyle\sum\limits_{k=0}^{\infty}}\frac{\left(it\right)^{k}}{k!}\left(\boldsymbol{P}_{l}\boldsymbol{J}\right)^{k}={\displaystyle\sum\limits_{k=0}^{\infty}}\frac{\left(it\right)^{k}}{k!}\boldsymbol{P}_{l}\boldsymbol{J}^{k}=\boldsymbol{P}_{l}{\displaystyle\sum\limits_{k=0}^{\infty}}\frac{\left(it\right)^{k}}{k!}\boldsymbol{J}^{k}=\boldsymbol{P}_{l}\mathrm{e}^{\mathrm{i}t\boldsymbol{J}}.$
@@
$\boldsymbol{P}_{l}\mathrm{e}^{t\boldsymbol{J}}=\mathrm{e}^{t\boldsymbol{P}_{l}\boldsymbol{J}}=\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}:\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right)\rightarrow\mathcal{D}_{l}\left(\mathbb{Z}_{p}\right).$
@@
$p_{r,v}(t)$ $=p^{l}{\displaystyle\int\limits_{v+p^{l}\mathbb{Z}_{p}}}Z_{0}\left(x,t\right)\ast\Omega\left(p^{l}\left|x-r\right|_{p}\right)dx=p^{l}{\displaystyle\int\limits_{v+p^{l}\mathbb{Z}_{p}}}\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)dx$
$={\displaystyle\int\limits_{\mathbb{Z}_{p}}}p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-v\right|_{p}\right)\mathrm{e}^{t\boldsymbol{J}^{\left(l\right)}}p^{\frac{l}{2}}\Omega\left(p^{l}\left|x-r\right|_{p}\right)dx,$
@@
13. Discussion
By $p$-adic QM, we mean quantum mechanics, in the sense of Dirac-von Neumann, with quantum states from a complex Hilbert space of the form $L^{2}(B)$, where $B$ is an open compact subset of $\mathbb{Q}_{p}^{N}$. In this framework, the time is a real variable, but the positions are vectors from $\mathbb{Q}_{p}^{N}$. The field of $p$-adic number $\mathbb{Q}_{p}$ is a paramount prototype of a “discrete space:” any continuous trajectory from $\mathbb{R}$ in to $\mathbb{Q}_{p}$ is constant. The motion in $\mathbb{Q}_{p}$ consists of a sequence of jumps. This assertion is also valid in any subset $B\subset\mathbb{Q}_{p}^{N}$; consequently, there are world lines in $B$, and so $p$-adic QM is not compatible with the theory of relativity.
$p$-Adic QM is a model of the Dirac-von Neumann formalism on a discrete space. Here, it is relevant to mention that there are several different types of $p$-adic QM. In some models, time is a $p$-adic variable; this type of model is outside the scope of the Dirac-von Neumann formalism. $p$-Adic QM with $p$-adic time was introduced in the 1980s by Vladimirov and Volovich [18]; the literature on it is extensive. Since the 1980s, a central problem has been to determine whether $p$-adic QM has a physical meaning.
In [39], the author and Mayes studied a $p$-adic version of the infinite potential well. This model describes the confinement of a particle in a $p$-adic ball. They rigorously solved the Cauchy problem for the Schrödinger equation and determined the stationary solutions. By dividing a $p$-adic ball into a finite number of sub-balls and using the wavefunctions of the infinite potential well, they constructed a continuous-time quantum walk (CTQW) on a fully connected graph, where each vertex corresponds to a sub-ball in the partition of the original ball. This approach shows that one can associate a CTQW with any $p$-adic Schrödinger equation. But this approach requires explicit formulas for solving the Cauchy problem for the Schrödinger equation, which is a very involved task and probably impossible for general equations, and, on top of that, the CTQWS produced do not include the ones used in quantum computing, [2]-[5]. Here, we recall that CTQWS are a particular case of continuous-time quantum Markov chains (CTQMCs). In the literature, CTQWs are usually constructed using adjacency matrices, whereas CTQMCs do not require them.
In [26], the author showed that a large class of $2$-adic Schrödinger equations is the scaling limit of CTQMCs. As a practical result, we construct new types CTQWs on graphs using two symmetric matrices. This construction includes, as a particular case, the CTQWs constructed using adjacency matrices. This paper did not include simulations of the CTQMCs, which are essential for understanding their dynamics.
The goal of this paper is to provide a draft of the theory of CTMCs associated with the $p$-adic Schrödinger equations, which contains as a particular case of the CTQWs constructed using adjacency matrices. Here, we study the CTMCs associated with certain $p$-adic heat equations, and the CTQMCs related to the Schrödinger equations obtained by Wick rotation. We solve the initial-value problems associated with the mentioned equations and provide numerical simulations for the corresponding CTQMCs. Behind the results presented in this paper is the conjecture, already formulated in [26], that general $p$-adic heat equations and the corresponding Schrödinger equations are scaling limits of CTMCs, respectively CTQMCs.
The connection between $p$-adic Schrödinger equations and CTQMCs demonstrates that $p$-adic QM has physical meaning. The $p$-adic Schrödinger operators are non-local, so from the beginning, ”spooky actions at a distance” are allowed. Then, interpreting the experimental confirmation of the violation of Bell’s inequalities as ”the universe is non-locally real,” which means that the need to choose between realism and non-locality implies that $p$-adic QM admits realism. The implications of this assertion are explored by the author in [43]; see also [24], [25] for the discussion of other $p$-adic models in QM.
Finally, it is worth noting that the developed framework remains valid if we replace the field of $p$-adic numbers with a more general ultrametric field, while keeping the state space as a Hilbert space of complex-valued functions. If the wavefunctions are not complex-valued, we cannot use the Dirac-von Neumann formalism. In particular, the study of CTQWs with $p$-adic transition probabilities is out of the scope of this work.